A normally distributed population has a mean of 1,214 and a standard deviation of 122. Find the probability that a single randomly selected element X of the population is between 1,100 and 1,300.


1
Expert's answer
2022-06-03T05:39:22-0400

"P(1,100<X<1,300)=P(z_1<z<z_2),"

where z1 and z2 are z-scores of 1,100 and 1,300.

Let's find them.

"z=\frac{x-\mu}{\sigma};" "\mu=1,214;" "\sigma=122."


"z_1=\frac{1,100-1,214}{122}\approx-0.93,"


"z_2=\frac{1,300-1,214}{122}\approx0.70."


So we have to rewrite the probability ang use z-table:

"P(1,100<X<1,300)=P(-0.93<z<0.70)="

"=0.7580-0.1762=0.5818=58.18\%."


Answer: 58.18%.